The minimum value of the function 2cos2x-cos4x in 0

Grade:12

2 Answers

Ramesh V
14 years ago

We have
f(x) = 2cos2x - cos4x
f ''(x) = -4sin2x + 4sin4x
= 4(sin4x - sin2x)
= 8cos3x.sin  
Now, 
f ''(x) = 0
8cos3xsinx= 0
cos3xsinx = 0
cos3x = 0 or sinx = 0
or x = 0, .
or x = 0, .   

(As 0  x  )
Now,
f(0) = 2cos0 - cos0 = 2 - 1 = 1  

f = 2cos- cos2= -2 - 1 = -3 
    
and f() = 2cos2- cos4= 2 - 1 = 1
Thus, the maximum value of f(x) is  and the minimum value of f(x) is -3.

--

Regards

Ramesh

Badiuddin askIITians.ismu Expert
14 years ago

Hi pallavi

 

2cos2x-cos4x

=2 cos2x -2cos22x +1

=-2(cos2x-1/2)2 +3/2

minimum value occur when cos2x-1/2 will be maximum

  =-2(-1-1/2)2 +3/2

  =-3


Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation. All the best pallavi.
 
 

Regards,

Askiitians Experts

Badiuddin